.60g of aluminum sulfate is produced when aluminum metal reacted with dilute sulfuric acid petrified is the law of conservation of mass is obeyed?
[Al=27,S=30, O=16, H=1].
1
Expert's answer
2018-08-15T04:12:27-0400
.60g of aluminum sulfate is produced when aluminum metal reacted with dilute sulfuric acid petrified is the law of conservation of mass is obeyed?
[Al=27,S=30, O=16, H=1].

Answer:
2Al + 3H2SO4 = Al2(SO4)3 + 3H2(g)
M=342 g/moll (Al2(SO4)3)
m1=60 g (Al2(SO4)3) = 60/342=0.1754 moll
m2= ? (Aluminum)
2 moll aluminum --------- 1 moll aluminum sulfate
X moll aluminum --------- 0.1754 moll aluminum sulfate
X= (0.1754 x 2) / 1
X= 0.3508 moll aluminum = 0.3508 x 27 = 9.4716
Y= 0.1754 x 3 x 98 = 51.5676 g dilute sulfuric acid
Z= 0.1754 x 3 x 2= 1.0524 g hydrogen

Law of conservation of mass:
m(Al) + m(H2SO4)=m(Al2(SO4)3) + m(H2)
9.4716 + 51.5676 = 61 (about)
60 + 1.0524 = 61 (about)